LET A and B be any two matrices such that B in the inverse of A.

1) Determine the relationship between the adjoint of A and the adjoint of B (5 MARKS)

2) Determine the relationship between the transpose of A and the transpose of B (5 marks).

Please assist.
1
Expert's answer
2020-04-22T17:42:48-0400

"B=A^{-1}"

Let "A^*" be adjoint of "A",

"B^*" be adjoint of "B".


1)

"A^{-1}=\frac{A^*}{detA}" and "B^{-1}=\frac{B^*}{detB}"

"B=A^{-1}=\frac{A^*}{detA}" (multyply by "B^{-1}" )

"BB^{-1}=\frac{A^*}{detA}B^{-1}"

"E=\frac{A^*}{detA}\frac{B*}{detB}"

"E=\frac{A^*}{detA}\frac{B*}{det(A^{-1})}"

"E=A^*B^*"

"B^*=(A^*)^{-1}"


2)

"A=B^{-1}"

"A^T=(B^{-1})^T"

"A^T=(B^T)^{-1}"

Another proof:

"AB=E"

"(AB)^T=E^T"

"B^TA^T=E"

"(B^T)^{-1}B^TA^T=(B^T)^{-1}E"

"A^T=(B^T)^{-1}"